OpenStudy (anonymous):

Verify the identity. tan x + pi /2= -cot x

OpenStudy (debbieg):

Surely it's tan (x + pi /2) = -cot x ??

OpenStudy (debbieg):

You can't just use the tan(a+b) identity because you would get some tan(pi/2) which is undefined. But you CAN use the definition of tana=sina/cosa. So you have: \[\Large \tan \left( x+\dfrac{\pi}{2} \right)=\frac{ \Large \sin \left( x+\dfrac{\pi}{2} \right) }{ \Large \cos \left( x+\dfrac{\pi}{2} \right) }\] Then just expand each of the num'r and den'r using the sum identities for sine and cosine, and I bet everything will fall into place. :)

OpenStudy (anonymous):

can you help me

OpenStudy (jdoe0001):

do you know the trig sum identities? what's sin(a + b) = ? what's cos(a + b) = ?

OpenStudy (anonymous):

sina+b+sinacosb

OpenStudy (anonymous):

?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

im lost

OpenStudy (jdoe0001):

hmm? well, where at?

OpenStudy (anonymous):

an you show me step by step

OpenStudy (jdoe0001):

ok, do you see the sum identities in the picture? what would be sin(a +b)? what would be cos(a+b)?

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

can you read this one => https://mathtestpreparation.com/Lessons/TrignSumADifForFig1.gif maybe

OpenStudy (jdoe0001):

so, what would be sin(a +b)? what would be cos(a+b)? as expanding the identity

OpenStudy (jdoe0001):

confused? where? :)

OpenStudy (anonymous):

how do you expand